a vending machine takes only nickels and dimes.there are 5 times as many dimes as nickels in the machine.the face value of the coins is $4.40.how many of each coin are in the machine?
Answer by checkley71(8403)

LET X=THE NUMBER OF NICKELS & 5X EQUAL THE NUMBER OF DIMES.
.05X+.10*5X=4.40
.05X+.50X=4.40
.55X=4.40
X=4.40/.55
X=8 IS THE NUMBER OF NICKELS.
5*8=40 IS THE NUMBER OF DIMES.
PROOF.
.05*8+.10*5*8=4.40
.40+4=4.40
4.40=4.40